OpenStudy (anonymous):

Please, almost done with math tonight...if 6, 29/2, and 23 are the first three terms of an arithmetic sequence, what is the sum of the first 18 terms of the sequence?

OpenStudy (anonymous):

you are all over the place with these

OpenStudy (anonymous):

common different is \(\frac{29}{2}-6=\frac{17}{2}\)

OpenStudy (anonymous):

i forget the formula for adding up the arithmetic sequence, but i bet we can find it

OpenStudy (anonymous):

ah it is \(\frac{n}{2}(a_1+a_n)\) in one version

OpenStudy (anonymous):

\[\frac{18}{2}(6+17(6+\frac{17}{2}))\]whatever that is

OpenStudy (anonymous):

I should send you a gift basket for all the help you've given me.